# GraphInsight: Unlocking Insights in Large Language Models for Graph Structure Understanding

### v2

@@ Abstract
Although Large Language Models (LLMs) have demonstrated potential in processing graphs, they struggle with comprehending graphical structure information through prompts of graph description sequences, especially as the graph size increases. We attribute this challenge to the uneven memory performance of LLMs across different positions in graph description sequences, known as “positional biases”. To address this, we propose GraphInsight, a novel framework aimed at improving LLMs’ comprehension of both macro- and micro-level graphical information. GraphInsight is grounded in two key strategies: 1) placing critical graphical information in positions where LLMs exhibit stronger memory performance, and 2) investigating a lightweight external knowledge base for regions with weaker memory performance, inspired by retrieval-augmented generation (RAG). Moreover, GraphInsight explores integrating these two strategies into LLM agent processes for composite graph tasks that require multi-step reasoning. Extensive empirical studies on benchmarks with a wide range of evaluation tasks show that GraphInsight significantly outperforms all other graph description methods (e.g., prompting techniques and reordering strategies) in understanding graph structures of varying sizes.
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in natural language processing (NLP) Shen et al. (2024); Naveed et al. (2023); Ge et al. (2023), enabling their initial applications across various data domains, such as graphs Chen et al. (2024); Wang et al. (2024c); Besta et al. (2024), time-series data Jin et al. (2023); Yu et al. (2023), tabular data Sui et al. (2024); Hegselmann et al. (2023), and other structured or semi-structured data types Ko et al. (2024); Perozzi et al. (2024). Among these domains, leveraging LLMs to tackle applications involving graphs has emerged as a burgeoning field of research, as graphs represent fundamental structures that capture intricate relationships and interactions in the real world Wang et al. (2021); Xu (2021). For example, Fatemi et al. have explored the potential of LLMs by converting various types of graphs, such as knowledge graphs Baek et al. (2023); Pan et al. (2024) and social network graphs Santra (2024); Babic (2023), into natural language descriptions, thereby enabling LLMs to perform question-answering tasks related to these graphs.
A key observation is that enhancing LLM performance in graph-related applications depends critically on LLMs’ ability to comprehend graph structures through natural language descriptions. Existing studies Shang & Huang (2024); Li et al. (2023) primarily utilizes two direct methods to transform graphs into text inputs for LLMs: the structural format transforming, such as adjacency matrices (termed as AM) or lists (termed as AL) and the sequential format transforming, such as edge-by-edge descriptions (termed as Raw Seq). However, extensive empirical studies Yuan et al. (2024) have shown that LLMs face significant challenges in understanding and reasoning about graph structures using current graph transformation methods, especially as graph size increases, leading to a “comprehension collapse”. As shown in Figure 1 (a), several common LLMs perform poorly on graph structure understanding tasks (see benchmarks in Section 5.1), and their comprehension declines sharply as the graph size increases, ultimately leading to complete failure.
@@
In this paper, we focus on enhancing the ability of LLMs to understand graph structures by developing new graph transforming methods and other novel techniques. To achieve this, we first analyze the inherent nature of graph understanding tasks in LLMs and the reasons behind the poor performance of existing methods. At its core, the challenge of LLM understanding graph structures can be viewed as a stringent long sequence comprehension problem Liu et al. (2024); Wang et al. (2024b). This challenge arises primarily due to two factors, as illustrated in Figure 1(b). First, graph structural information is conveyed to LLMs solely through language descriptions. As the graph size increases, there descriptions becomes longer, challenging LLMs’ ability to comprehend long-sequence inputs. Second, to accurately understand the graph’s structure, LLMs must retain every detail in the natural language description. Any memory lapse, especially regarding critical nodes or edges (e.g., central nodes, bridging edges), notably impairs the LLMs’ ability to infer the correct structure from the description.
@@
However, the challenge is further exacerbated by the uneven memory performance of LLMs across different positions within long sequences—a phenomenon known as “positional biases” Tang et al. (2023); Hsieh et al. (2024); Zhang et al. (2024), which mainly stems from limitations in the attention mechanism Xiao et al. (2024) and the internal memory capacity An et al. (2024) of LLMs. For example, as shown in Figure 2(a), many studies indicate that LLMs generally perform better at understanding the head and tail of sequences (i.e., strong memory regions), with a noticeable dip in performance in middle parts (i.e., weak memory regions), termed as “lost-in-the-middle” Liu et al. (2024). Consequently, LLMs struggle to meet the stringent requirements for understanding graph structures from descriptive sequences, due to they always lose some positions in the sequence. This issue becomes particularly pronounced as the graph size increases (i.e., sequence lengthens), leading to “comprehension collapse,” as mentioned in Figure 1(a).
Figure 2: Analysis on Positional Biases of LLMs
@@
Thus, improving LLMs’ ability to comprehend graph structures relies on improving their ability to retain information across different sequence positions in graph descriptions, thereby reducing the impact of positional biases. To this end, our work is grounded in two high-level insights: 1) Taking advantage of the Strong Memory Regions: For sequence positions where LLMs demonstrate strong memory capabilities, we strategically place descriptions of critical nodes or edges in these locations. 2) Compensating for the Weak Memory Regions: we construct a lightweight external knowledge base for positions where LLMs show weak memory performance, inspired by the concept of retrieval-augmented generation (RAG). As illustrated in Figure 1(b), following the above “different horses for different courses” principle, the LLMs’ comprehension of graph description sequences improves from Curve $A$ to Curve $B$.
@@ 2.1 LLMs’ Understanding of Graph Structures Through Description Sequences
Research on the ability of LLMs to understand graph structures by inputting natural language descriptions of graphs into these models has become an emerging area of study. Existing work has primarily focused on designing evaluation benchmarks for graph structure understanding tasks and analyzing the results. For example, NLGraph and GPT4Graph provided preliminary empirical assessments of LLMs’ performance on coarse-grained graph structure tasks, indicating that this research is still in its early stages. Subsequently, benchmarks such as GraphEval2000 Wu et al. (2024b), GraphArena Tang et al. (2024), GraphInstruct Luo et al. (2024), and GraCoRe Yuan et al. (2024) have concentrated on fine-grained tasks, such as graph theory problems. Ge et al. also explored the impact of the sequence of graph descriptions on LLMs comprehension, though their study was limited to a few tasks and lacked deep analysis. Unlike previous work, our framework aims to enhance LLMs’ capabilities in graph structure understanding tasks.
@@ 2.2 Positional Bias in Long Sequences for LLMs
Some recent works have attempted to mitigate the positional bias issue in LLMs. For instance, Zhang et al. and Hsieh et al. introduced attention instructions to guide the LLMs to focus on specific sequence segments. An et al. proposed a training solution using a synthesized long- sequence question-answer dataset to alleviate the “lost-in-the-middle” problem. Additionally, Wu et al. suggested interpolating positional encodings by manipulating position indices, a method that requires fine-tuning within the pre-trained context window. Note that these broad strategies are not specifically designed for graphs and do not directly apply to graph description sequences; instead, they are generally implemented during the training/fine-tuning phases of LLMs. Therefore, these methods are orthogonal to our framework and could further enhance its performance if computational resources permit.
@@ 3 Preliminaries
In this section, we define the basic format of graph description sequences and outline the capability of LLMs on these sequences. We then introduce two levels of graph understanding tasks, laying the foundation for our framework.
@@
The description $\mathcal{T}$ consists of the descriptions of all edges $e_{ij}\in E$.
Based on Definition 1, the LLMs’ ability to understand $\mathcal{T}$ can be quantified as shown in Definition 2.
@@
Here, $\Psi(p)$ denotes the positional bias curve, which represents the distribution of an LLMs’ inherent comprehension at position $p$. As mentioned in Figure 2, $\Psi(p)$ typically displays stronger comprehension at the head and tail of a sequence, with weaker comprehension in the middle, forming a U-shape curve. In this paper, we follow above assumption and deine the positions corresponding to the head $\alpha\%$ and tail $\beta\%$ of the sequence as strong memory regio ns, and the rest as weak memory regions.
Next, we introduce two levels of graph understafnding tasks, as shown in Definition 3.
@@
Integrating on existing benchmarks Wang et al. (2024a); Yuan et al. (2024), we categorize LLM graph structure understanding into two levels: 1) Macro-level tasks, involving coarse-grained reasoning related to the overall graph structure (e.g., node counting, connectivity detection, cycle detection); 2) Micro-level tasks, focusing on fine-grained reasoning related to local structures (e.g., direct connection detection, node degree calculation, leaf node identification, neighbor recognition). Moreover, some composite graph understanding tasks may require multiple micro-level reasoning steps (e.g., complete subgraph verification, third-order neighbor identification).
@@ 4 Methodology
Figure 3: Framework of GraphInght.
@@ 4.2 Importance-based Macro-level Graph Understanding
For macro-level graph understanding tasks, guiding LLMs to focus on the memory and comprehension of key elements within the graph structure can potentially enhance their reasoning performance. Thus, an intuitive idea is to align the LLMs’ comprehension distribution across different sequence positions (i.e., $\Psi(p)$) with the importance distribution of these positions in the graph description sequence. Let $\Phi(p)$ represent the importance of each corresponding position $p$ in the sequence $\mathcal{T}$ relative to the graph structure. In this context, our objective is:
@@
Specifically, for the graph description decomposition, given a graph $G$ and its corresponding graph description sequence $\mathcal{T}$, we first iteratively calculate the PageRank score $\text{PR}(v)$ [^2] PR can be replaced with others describe the importance of graph structures. Designing graph structure importance remains an open problem Besta et al. (2023); Liu & Gao (2023); Geng et al. (2022), and is orthogonal to our work. for each node $v\in V$ in $G$ as follows:
@@
After calculating the PageRank scores, we sort the nodes in descending order. Starting from the node with the highest score, we iteratively construct subgraphs centered on each node, including directly connected neighbors with lower degrees, provided the connecting edges haven’t been used in other subgraphs. This ensures that each edge is only included in one subgraph, which guarantees that the total length of the merged subgraph descriptions equals the original graph description sequence. This is because, as per Definition 1, we describe the graph edge by edge. The descriptions for these subgraphs $G_{1},G_{2},\dots,G_{k}$ are denoted as $t_{1},t_{2},\dots,t_{k}$. Then, each subgraph description $t_{i}$ is formed by combining the descriptions of all edges within the subgraph $G_{i}$, according to the graph description defined previously in Definition 1. The importance $\mathcal{I}(t_{i})$ of each subgraph description $t_{i}$ is defined as the PageRank score of its central node $v_{c_{i}}$: $\mathcal{I}(t_{i})=\text{PR}(v_{c_{i}})$.
Finally, the subgraph descriptions are reordered based on their importance and organized within the strong memory regions of the LLM, as defined in Definition 2. Specifically, the head $\alpha\%$ and the tail $\beta\%$ of the graph description sequence are designated as the strong memory regions. Consequently, the most important subgraph descriptions are prioritized and placed in these regions, while the remaining descriptions, sorted by importance, occupy the middle, considered the weak memory region. The final graph description sequence $\mathcal{\hat{T}}$ is:
@@ 4.3 RAG-based Micro-level Graph Understanding
Next, we introduce the construction of our framework’s RAG knowledge base, called the “GraphRAG base”, and the corresponding RAG process, termed the “GraphRAG process”. Note that existing RAG techniques Ghosh et al. (2024); Rorseth et al. (2024); Sojitra et al. (2024); Das et al. (2024) and optimizations are orthogonal to our framework and can further enhance RAG quality, but here we focus only on the most basic RAG methods.
@@
Conventional RAG base typically require substantial storage and rely on extensive structured or unstructured data (e.g., documents, knowledge graphs). In contrast, our GraphRAG base, denoted as $\mathcal{K}$, demands minimal storage overhead. Specifically, for the graph description sequence $\mathcal{\hat{T}}$ generated by the importance-based description reorganization, the nodes and edges of the subgraph structures corresponding to the weak memory regions within $\mathcal{\hat{T}}$ will be stored as the GraphRAG base. The proportion of stored subgraph structures, denoted as $\gamma\%$, is adjustable. Since the subgraph structures corresponding to the weak memory regions have already been ranked by importance, we can conveniently select and store only the top $\gamma\%$ of these structures. Moreover, to facilitate efficient RAG retrieval, we store the node and edge information of the subgraph structures that need to be included in the GraphRAG base separately: $\mathcal{K}=\{\mathcal{K}_{node},\mathcal{K}_{edge}\}$. Among them,
@@
Based on the GraphRAG base $\mathcal{K}$, when augmented graph information is needed for a micro-level graph understanding task $q$, we first extract/identify all nodes from $q$ by their names or identifiers. This can be accomplished using various entity recognition Li et al. (2022) or semantic parsing models Lewis et al. (2020); Zhu et al. (2024):
$V_{q}=\text{EntityRecognition}(q)$ (7)
where $V_{q}$ represents the set of node entities extracted from the task $q$. Next, based on the $V_{q}$, we perform retrieval operations on $\mathcal{K}_{edge}$. Specifically, for $\mathcal{K}_{\text{node}}$, we directly retrieve each node and its degree information based on the nodes in $V_{q}$, resulting in the augmented information set $\mathcal{K}_{\text{node}}^{q}$:
@@
Our GraphInsight framework incorporates two key techniques that can be seamlessly integrated into the agent-based processes of LLMs to enhance the performance of such multi-step reasoning tasks:

### v3

@@ Abstract
Although Large Language Models (LLMs) have demonstrated potential in processing graphs, they struggle with comprehending graphical structure information through prompts of graph description sequences, especially as the graph size increases. We attribute this challenge to the uneven memory performance of LLMs across different positions in graph description sequences, known as “Positional bias”. To address this, we propose GraphInsight, a novel framework aimed at improving LLMs’ comprehension of both macro- and micro-level graphical information. GraphInsight is grounded in two key strategies: 1) placing critical graphical information in positions where LLMs exhibit stronger memory performance, and 2) investigating a lightweight external knowledge base for regions with weaker memory performance, inspired by retrieval-augmented generation (RAG). Moreover, GraphInsight explores integrating these two strategies into LLM agent processes for composite graph tasks that require multi-step reasoning. Extensive empirical studies on benchmarks with a wide range of evaluation tasks show that GraphInsight significantly outperforms all other graph description methods (e.g., prompting techniques and reordering strategies) in understanding graph structures of varying sizes.
GraphInsight: Unlocking Insights in Large Language Models
for Graph Structure Understanding
Yukun Cao1,2thanks: These authors contributed equally to this work.  Shuo Han1,211footnotemark: 1  Zengyi Gao1,2 Zezhong Ding1,2Xike Xie1,2thanks: Corresponding authorS. Kevin Zhou1,3 1 University of Science and Technology of China, China 2 Data Darkness Lab, MIRACLE Center, USTC, China 3 MIRACLE Center, USTC, China {ykcho, shuo.han, gzy02, zezhongding}@mail.ustc.edu.cn {xkxie, skevinzhou}@ustc.edu.cn
@@ 1 Introduction
Large language models (LLMs) have demonstrated remarkable capabilities in natural language processing (NLP) Shen et al. (2024); Naveed et al. (2023); Ge et al. (2023), enabling their initial applications across various data domains, such as graphs Chen et al. (2024b); Wang et al. (2024c); Besta et al. (2024), time-series data Jin et al. (2023); Yu et al. (2023), tabular data Sui et al. (2024); Hegselmann et al. (2023), and other structured or semi-structured data types Ko et al. (2024); Perozzi et al. (2024). Among these domains, leveraging LLMs to tackle applications involving graphs has emerged as a burgeoning field of research, as graphs represent fundamental structures that capture intricate relationships and interactions in the real world Wang et al. (2021); Xu (2021). For example, Fatemi et al. have explored the potential of LLMs by converting various types of graphs, such as knowledge graphs Baek et al. (2023); Pan et al. (2024) and social network graphs Santra (2024); Babic (2023), into natural language descriptions, thereby enabling LLMs to perform question-answering tasks related to these graphs.
@@
A key observation is that enhancing LLM performance in graph-related applications depends critically on LLMs’ ability to comprehend graph structures through natural language descriptions. Existing studies Shang and Huang (2024); Li et al. (2023) primarily utilize two direct methods to transform graphs into text inputs for LLMs: the structural format transforming, such as adjacency matrices (termed as AM) or lists (termed as AL) and the sequential format transforming, such as edge-by-edge descriptions (termed as Raw Seq). However, extensive empirical studies Yuan et al. (2024) have shown that LLMs face significant challenges in understanding and reasoning about graph structures using current graph transformation methods, especially as graph size increases, leading to a “comprehension collapse”. As shown in Figure 1 (a), several common LLMs perform poorly on graph structure understanding tasks (see benchmarks in Section 5.1), and their comprehension declines sharply as the graph size increases, ultimately leading to complete failure.
@@
In this paper, we focus on enhancing the ability of LLMs to understand graph structures by developing new graph-transforming methods and other novel techniques. To achieve this, we first analyze the inherent nature of graph understanding tasks in LLMs and the reasons behind the poor performance of existing methods. At its core, the challenge of LLM understanding graph structures can be viewed as a stringent long sequence comprehension problem Liu et al. (2024); Wang et al. (2024b). This challenge arises primarily due to two factors, as illustrated in Figure 1(b). First, graph structural information is conveyed to LLMs solely through language descriptions. As the graph size increases, their descriptions become longer, challenging LLMs’ ability to comprehend long-sequence inputs. Second, to accurately understand the graph’s structure, LLMs must retain every detail in the natural language description. Any memory lapse, especially regarding critical nodes or edges (e.g., central nodes, bridging edges), notably impairs the LLMs’ ability to infer the correct structure from the description.
@@
However, the challenge is further exacerbated by the uneven memory performance of LLMs across different positions within long sequences—a phenomenon known as “positional bias” Tang et al. (2023); Hsieh et al. (2024); Zhang et al. (2024a), which mainly stems from limitations in the attention mechanism Xiao et al. (2024) and the internal memory capacity An et al. (2024) of LLMs. For example, as shown in Figure 2(a), many studies indicate that LLMs generally perform better at understanding the head and tail of sequences (i.e., strong memory regions), with a noticeable dip in performance in middle parts (i.e., weak memory regions), termed as “lost-in-the-middle” Liu et al. (2024). Consequently, LLMs struggle to meet the stringent requirements for understanding graph structures from descriptive sequences, due to they always lose some positions in the sequence. This issue becomes particularly pronounced as the graph size increases (i.e., sequence lengthens), leading to “comprehension collapse,” as mentioned in Figure 1(a).
Figure 2: Analysis on Positional bias of LLMs
@@
Thus, improving LLMs’ ability to comprehend graph structures relies on improving their ability to retain information across different sequence positions in graph descriptions, thereby reducing the impact of positional bias. To this end, our work is grounded in two high-level insights: 1) Taking advantage of the Strong Memory Regions: For sequence positions where LLMs demonstrate strong memory capabilities, we strategically place descriptions of critical nodes or edges in these locations. 2) Compensating for the Weak Memory Regions: we construct a lightweight external knowledge base for positions where LLMs show weak memory performance, inspired by the concept of retrieval-augmented generation (RAG). As illustrated in Figure 1(b), following the above “different horses for different courses” principle, the LLMs’ comprehension of graph description sequences improves from Curve $A$ to Curve $B$.
@@ 2.1 LLMs’ Understanding of Graph Structures Through Description Sequences
Research on the ability of LLMs to understand graph structures by inputting natural language descriptions of graphs into these models has become an emerging area of study. Existing research has focused on developing evaluation benchmarks for graph structure understanding tasks and analyzing their results. For instance, Wang et al. and Guo et al. conducted preliminary empirical assessments of LLMs on coarse-grained graph structure tasks, indicating the nascent stage of this field. Subsequent benchmarks, such as GraphEval2000 Wu et al. (2024b), GraphArena Tang et al. (2024), GraphInstruct Luo et al. (2024), and GraCoRe Yuan et al. (2024), have shifted focus to fine-grained tasks, including graph theory problems. Ge et al. also explored the impact of the sequence of graph descriptions on LLMs comprehension, though their study was limited to a few tasks and lacked deep analysis. Unlike previous work, GraphInsight aims to enhance LLMs’ capabilities in graph structure understanding tasks.
@@ 2.2 Positional Bias in Long Sequences for LLMs
Some recent works have attempted to mitigate the positional bias issue in LLMs. For instance, Zhang et al.; Hsieh et al. introduced attention instructions to guide the LLMs to focus on specific sequence segments. An et al. proposed training with a synthesized long-sequence QA dataset to mitigate the "lost-in-the-middle" problem. Wu et al. proposed interpolating positional encodings via index adjustment, requiring fine-tuning within the pre-trained context window. Note that these broad strategies are not specifically designed for graphs and do not directly apply to graph description sequences; instead, they are generally implemented during the training/fine-tuning phases of LLMs. Therefore, these methods are orthogonal to GraphInsight and could further enhance its performance if computational resources permit.
@@ 3 Preliminaries
In this section, we define the basic format of graph description sequences, outline our assessment of LLM capabilities on these sequences, and introduce two levels of graph understanding tasks as the foundation of our framework.
@@
The description $\mathcal{T}$ consists of the descriptions of all edges $e_{ij}\in E$. Based on Definition 1, the LLMs’ ability to understand $\mathcal{T}$ can be quantified as shown in Definition 2.
@@
Here, $\Psi(p)$ denotes the positional bias curve, which represents the distribution of an LLM’s inherent comprehension at position $p$. As mentioned in Figure 2, $\Psi(p)$ typically displays stronger comprehension at the head and tail of a sequence, with weaker comprehension in the middle, forming a U-shape curve. In this paper, we follow the above assumption and define the positions corresponding to the head $\alpha\%$ and tail $\beta\%$ of the sequence as strong memory regions, and the rest as weak memory regions.
Next, we introduce two levels of graph understanding tasks, as shown in Definition 3.
@@
Integrating on existing benchmarks Wang et al. (2024a); Yuan et al. (2024), we categorize LLM graph structure understanding into two levels: 1) Macro-level, involving coarse-grained reasoning related to the overall graph structure (e.g., node counting, connectivity detection, cycle detection); 2) Micro-level, focusing on fine-grained reasoning related to local structures (e.g., direct connection detection, node degree calculation, leaf node identification, neighbor recognition). Moreover, some composite graph understanding tasks may require multiple micro-level reasoning steps (e.g., complete subgraph verification, and third-order neighbor identification).
@@ 4 Methodology
Figure 3: Framework of GraphInsight.
@@ 4.2 Importance-based Macro-level Graph Understanding
For macro-level graph understanding tasks, guiding LLMs to focus on the memory and comprehension of key elements within the graph structure can potentially enhance their reasoning performance. Thus, the intuitive idea is to align the LLMs’ comprehension distribution across different sequence positions (i.e., $\Psi(p)$) with the importance distribution of these positions in the graph description sequence. Let $\Phi(p)$ represent the importance of each corresponding position $p$ in the sequence $\mathcal{T}$ relative to the graph structure. In this context, our objective is:
@@
Specifically, for the graph description decomposition, given a graph $G$ and its corresponding graph description sequence $\mathcal{T}$, we first iteratively calculate the PageRank score $\text{PR}(v)$ [^2] PR can be replaced with others describe the importance of graph structures. Designing graph structure importance remains an open problem Besta et al. (2023); Liu and Gao (2023); Geng et al. (2022), and is orthogonal to our work. for each node $v\in V$ in $G$ as follows:
@@
After calculating the PageRank scores, we sort the nodes in descending order. Starting from the node with the highest score, we iteratively construct subgraphs centered on each node, including directly connected neighbors with lower degrees, provided the connecting edges haven’t been used in other subgraphs. This ensures that each edge is only included in one subgraph, guaranteeing that the total length of the merged subgraph descriptions equals the original graph description sequence. This is because, as per Definition 1, we describe the graph edge by edge. The descriptions for these subgraphs $G_{1},G_{2},\dots,G_{k}$ are denoted as $t_{1},t_{2},\dots,t_{k}$. Then, each subgraph description $t_{i}$ is formed by combining the descriptions of all edges within the subgraph $G_{i}$, according to the graph description defined previously in Definition 1. The importance $\mathcal{I}(t_{i})$ of each subgraph description $t_{i}$ is defined as the PageRank score of its central node $v_{c_{i}}$: $\mathcal{I}(t_{i})=\text{PR}(v_{c_{i}})$.
Finally, the subgraph descriptions are reordered based on their importance and organized within the strong memory regions of the LLM, as defined in Definition 2. Specifically, the head $\alpha\%$ and the tail $\beta\%$ of the graph description sequence are designated as the strong memory regions. Consequently, the most important subgraph descriptions are prioritized and placed in these regions, while the remaining descriptions, sorted by importance, occupy the middle and are considered the weak memory region. The final graph description sequence $\mathcal{\hat{T}}$ is:
@@ 4.3 RAG-based Micro-level Graph Understanding
Next, we introduce the construction of our framework’s RAG knowledge base, termed “GraphRAG base", and the corresponding RAG process, termed the “GraphRAG process". Note that existing RAG techniques Ghosh et al. (2024); Rorseth et al. (2024); Sojitra et al. (2024); Das et al. (2024) and optimizations are orthogonal to our framework and can further enhance RAG quality, but here we focus only on the most basic RAG methods.
@@
Conventional RAG base typically require substantial storage and rely on extensive structured or unstructured data (e.g., documents, knowledge graphs). In contrast, our GraphRAG base, denoted as $\mathcal{K}$, demands minimal storage overhead. Specifically, for the graph description sequence $\mathcal{\hat{T}}$ generated by the importance-based description reorganization, the nodes and edges of the subgraph structures corresponding to the weak memory regions within $\mathcal{\hat{T}}$ will be stored as the GraphRAG base. The proportion of stored subgraph structures denoted as $\gamma\%$, is adjustable. Since the subgraph structures corresponding to the weak memory regions have already been ranked by importance, we can conveniently select and store only the top $\gamma\%$ of these structures. Moreover, to facilitate efficient RAG retrieval, we store the node and edge information of the subgraph structures that need to be included in the GraphRAG base separately: $\mathcal{K}=\{\mathcal{K}_{node},\mathcal{K}_{edge}\}$. Among them,
@@
Based on the GraphRAG base $\mathcal{K}$, when augmented graph information is needed for a micro-level graph understanding task $q$, we first extract/identify all nodes from $q$ by their names or identifiers. This can be accomplished using various entity recognition Li et al. (2022) or semantic parsing models Lewis et al. (2020); Zhu et al. (2024): $V_{q}=\text{EntityRecognition}(q)$, where $V_{q}$ represents the set of node entities extracted from the task $q$. Next, based on the $V_{q}$, we perform retrieval operations on $\mathcal{K}_{edge}$. Specifically, for $\mathcal{K}_{\text{node}}$, we directly retrieve each node and its degree information based on the nodes in $V_{q}$, resulting in the augmented information set $\mathcal{K}_{\text{node}}^{q}$:
@@
GraphInsight framework incorporates two key techniques that can be seamlessly integrated into the agent-based processes of LLMs to enhance the performance of such multi-step reasoning tasks:
  • $\bullet$
@@
Models Tasks Raw Seq. GraphToken Prompting Methods Reordering Methods Structural Methods GraphInsight
COT FS BAG BFS DFS SP AL AM
Mistral-7B Overall 0.5476 0.2644 0.5315 0.4671 0.5145 0.5481 0.5482 0.5486 0.5320 0.1922 0.6811
Macro 0.5222 0.3342 0.4323 0.5439 0.4249 0.5212 0.5205 0.5222 0.5001 0.1808 0.5846
Micro 0.5635 0.2208 0.5935 0.4192 0.5705 0.5650 0.5655 0.5651 0.5519 0.1994 0.7425
Llama-3-8B Overall 0.4513 0.3480 0.4504 0.4173 0.4358 0.4551 0.4591 0.4559 0.4048 0.1434 0.6765
Macro 0.4379 0.3989 0.4187 0.3702 0.4060 0.4381 0.4376 0.4404 0.4336 0.1198 0.5422
Micro 0.4597 0.3161 0.4702 0.4468 0.4545 0.4658 0.4726 0.4656 0.3869 0.1581 0.7605
Qwen2-7B Overall 0.5640 0.3550 0.5663 0.5400 0.5677 0.5381 0.5290 0.5292 0.5595 0.2129 0.7695
Macro 0.5644 0.4493 0.5423 0.5366 0.6447 0.6017 0.5913 0.5958 0.5223 0.1915 0.6587
Micro 0.5637 0.2960 0.5813 0.5421 0.5196 0.4984 0.4901 0.4876 0.5828 0.2262 0.8387
Llama-3-8B-262k Overall 0.7183 0.2703 0.5279 0.5825 0.6231 0.7150 0.7166 0.7147 0.6603 0.5038 0.8285
Macro 0.6218 0.4198 0.5751 0.4652 0.5115 0.6189 0.6221 0.6183 0.5292 0.3529 0.6928
Micro 0.7786 0.1770 0.4985 0.6559 0.6928 0.7752 0.7757 0.7750 0.7422 0.5982 0.9401
Vicuna-7B Overall 0.1273 0.1794 0.1147 0.1041 0.1066 0.1276 0.1277 0.1277 0.0862 0.0345 0.3557
Macro 0.1267 0.1575 0.1164 0.0877 0.0995 0.1266 0.1259 0.1263 0.0567 0.0409 0.2296
Micro 0.1276 0.2016 0.1137 0.1144 0.1111 0.1282 0.1289 0.1285 0.1046 0.0306 0.4346